# ReinWiFi: A Reinforcement-Learning-Based Framework for the Application-Layer QoS Optimization of WiFi Networks

### v1

@@ V-B Online Q-network Training
$\mathcal{A}_{t}^{i}=\begin{cases}\operatorname*{arg\,min}Q^{i}(\hat{\mathcal{S}}_{t},\mathcal{A}^{i};\boldsymbol{\theta}_{t})&\text{with probability }1-\epsilon_{t},\\ \mathcal{A}^{i}\sim\mathrm{Unif}(\mathscr{A}^{i})&\text{with probability }\epsilon_{t},\end{cases}$ (19)
@@ VI Experiments
The proposed ReinWiFi system is implemented in a WiFi network with one HONOR XD30 AP and $3$ UEs each equipped with a TP-Link TL-WDN6200 USB WiFi adapter in the experiment [^1] The source code of implementation is available online in https://github.com/QianrenLi/ReinWiFi.. Denote the AP as $u_{0}$ and the three UEs as $u_{1},u_{2},u_{3}$, respectively. The network is working on the $5$G WiFi band following the IEEE 802.11ac specification. The real-time controller is implemented in a laptop with Intel Core i7-8750H CPU and Ubuntu 20.04 operating system. An Ethernet connection with a maximum data rate of $1$ Gbps is employed to facilitate communication between the controller and the AP. Moreover, we implement a Linux module to adapt the CW sizes of TL-WDN6200 adapters in real-time from user space. Hence, the controller can collect the QoS observations from UEs and notify the scheduling actions via WiFi, such that the UEs’ transmission scheduling can be adjusted accordingly.
@@
In the experiment, the duration of the scheduling period is 1 second, the CW size takes values from $\{2^{i}-1\mid i=1,2,\ldots,10\}$, and throughput limitation takes values from $\{\frac{i}{20}r_{{i},{j}}^{{m},\max}\mid i=0,1,\ldots,20\}$, where $r_{{i},{j}}^{{m},\max}$ = 600 Mbps. Moreover, in addition to the background interference, the interfering traffic between two interfering UEs, denoted as $u_{4}$ and $u_{5}$, is generated with a random data rate and BE priority in the same channel.
@@
The performance comparison of the proposed framework and the two baselines in the first $5$ test scenarios is illustrated in Fig. 1, where the online training of the Q-network is not applied in the proposed framework and the Baseline 2. It can be observed that the proposed Q-network offline trained via imitators significantly outperforms the conventional EDCA mechanism. Moreover, the performance gain of the Baseline 2 over Baseline 1 demonstrates the necessity of the throughput limitation, which has never been investigated in the existing literature.

### v2

@@ V-B Online Q-network Training
$\mathcal{A}_{t}^{i}=\begin{cases}\operatorname*{arg\,min}Q^{i}(\hat{\mathcal{S}}_{t},\mathcal{A}^{i};\boldsymbol{\theta}_{t})&\text{w.p. }1-\epsilon_{t},\\ \mathcal{A}^{i}\sim\mathrm{Unif}(\mathscr{A}^{i})&\text{w.p. }\epsilon_{t},\end{cases}$ (19)
@@ VI Experiments
Figure 1: Illustration of testbed.
As illustrated in Fig. 1 The proposed ReinWiFi system is implemented in a WiFi network with one HONOR XD30 AP and $3$ UEs each equipped with a TP-Link TL-WDN6200 USB WiFi adapter in the experiment [^1] The demo video is available in http://lasso.eee.sustech.edu.cn/reinwifi/; The source code is available in https://github.com/QianrenLi/ReinWiFi.. Denote the AP as $u_{0}$ and the three UEs as $u_{1},u_{2},u_{3}$, respectively. The network is working on the $5$G WiFi band following the IEEE 802.11ac specification. The real-time controller is implemented in a laptop with Intel Core i7-8750H CPU and Ubuntu 20.04 operating system. An Ethernet connection with a $1$ Gbps data rate is employed to facilitate communication between the controller and the AP. Moreover, we implement a Linux module to adapt the CW sizes of adapters in real-time from user space. The controller can collect the QoS observations from UEs and notify the scheduling actions via WiFi, such that the UEs’ transmission scheduling can be adjusted accordingly.
@@
In the experiment, the scheduling period duration is 1 second, the CW size takes values from $\{2^{i}-1\mid i=1,2,\ldots,10\}$, and throughput limitation takes values from $\{\frac{i}{20}r_{{i},{j}}^{{m},\max}\mid i=0,1,\ldots,20\}$, where $r_{{i},{j}}^{{m},\max}$ = 600 Mbps. Moreover, in addition to the background interference, the interfering traffic between two interference UEs, namely $u_{4}$ and $u_{5}$, is generated with a random data rate and BE priority in the same channel.
@@
The performance comparison of the proposed framework and the two baselines in the first $5$ test scenarios is illustrated in Fig. 2, where the online training is not applied in the proposed framework and the Baseline 2. It can be observed that the proposed Q-network offline trained via imitators significantly outperforms the conventional EDCA mechanism. Moreover, the performance gain of the Baseline 2 over Baseline 1 demonstrates the necessity of the throughput limitation, which has never been investigated in the existing literature.